# Hyperdimensional Computing for Node Classification and Link Prediction

### v2

@@ 3.4 Representational Power of GNN
Several authors have investigated the representational power of GNN, including their ability to distinguish between different graph structures [59], e.g., using the Weisfeiler-Lehman (WL) graph isomorphism test [56, 19] , or to learn a set of functions on graphs [44], or the equivalence of the preceding two notions of expressivity[8]. The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors [56, 2]. The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors [19]. Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into distinct encodings. As shown by [59, 33], GNN are at most as powerful as the WL test in distinguishing graph structures if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme: $\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$ where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations [59].
3.5 Limitations of Previous HD Graph Methods
The HD Computing based graph-level learning methods introduced by Nunes et al. [38] and Poduval et al. [43] are designed for similarity-based graph retrieval and graph classification, as opposed to node classification or link prediction. Nunes et al. [38] constructs the Graph HD vector using node features derived using PageRank [5]. Poduval et al. [43] assign random HD-vectors to nodes and considers them as node features to build a Graph HD vector. Thus, both methods ignore node features present in the graph.
RelHD [24] constructs node representations for node classification tasks and uses the resulting representation to design a novel hardware chip design with a scalable processing in-memory (PIM) architecture based FeFET. However, RelHD has a significant limitation with respect how it encodes the node features into HD vectors. Specifically, it treats node features as dimension-wise indicator variables and computes the HD vector representation of each node by first generating $d$ random basis vector. Let $\Upsilon_{p}$ be such a random basis Hyper vector. Random basis vectors $\Upsilon_{p}$ are bundled/accumulated if feature at index $p$ is present i.e HD vector for a node $u$ is computed using basis vectors and raw features $\mathbf{x}_{u}=[x_{u1},\ldots,x_{ud}]$ using the following formula:- $\oplus_{p\in\{1,\ldots,d\}}(\mathbbm{1}(x_{up}\neq 0)\cdot\Upsilon_{p})$. Consequently, this method is not suitable for graphs with continuous or integer-valued node attributes in continuous form or even simple bag-of-words counts. Note that because HyperNode [31] also utilizes the same mechanism for constructing HD vector encodings of the nodes, it suffers from the same limitations as RelHD.
Furthermore, none of the aforementioned methods [38, 43, 24, 31] support link prediction on graphs.
4 Hyper-Dimensional Graph Learner
In essence, our design of HDGL aims to achieve a high level of expressive power similar to graph neural networks (GNN). We utilize HD computing operations on node representations, taking advantage of HD operations—Bundling ($\oplus$), Binding ($\otimes$), and Rotation ($\Pi$) —and their properties. This ensures that HDGL, akin to GNN, not only aspires to build expressive power but also learns to map similar graphs to similar representations.
The proposed HD learner carries out 3 major steps: (i). Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$. (ii). Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$. (iii). Use $\mathbf{z}_{v}$ to predict $y_{v}$ or predict links/edges in the graph. We proceed to describe each of these steps in detail.
Figure 1: An example mapping node features to a binary space using Random Hyperplanes. A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side. Utilizing multiple hyperplanes results in a binary sketch for each point.
4.1 Mapping Node Features to HD-space
@@
The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$. We do this using random hyperplane tessellations [12], a type of Locality Sensitive Hashing [21, 7]. The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane. With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node. More precisely, let $\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{a}_{1}^{T},\cdots\mathbf{a}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$. For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$. Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$. The primary purpose of adding the $\Gamma$ offset is to ensure that the not all sampled hyperplanes pass through the origin. When dealing with evenly spread points, typically in the context of real-valued vectors, setting $\Gamma$ to zero is often appropriate.
Note that the resulting binary sketch of each node is in fact a $\beta$-dimensional binary HD vector. Nodes with similar features will have similar HD representations; and nodes with dissimilar features will have dissimilar HD representations.
4.2 Neighborhood Aggregation Scheme
@@
Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows: $\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}$ where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
@@
Bundling for generalizing across similar nodes. We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets). Hence, we propose to use Bundling to realize $\varphi_{k}$. More precisely, $\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$. This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective. However, as noted by [59], GNNs that use the mean to aggregate features over node neighborhoods [26, 18] perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
It is worth noting that the Binding operation $\otimes$ would be inappropriate in this context because binding of two similar vectors results in dissimilar vectors. Thus, if Binding were used to realize $\varphi_{k}$, similar yet distinct local neighborhoods structures would yield different HD representations, adversely impacting HDGL’s ability to generalize across such graph structures. To complicate matters, Binding is also invertible i.e $\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$; and hence, it could discard information that is useful for downstream task e.g., node label prediction.
Binding for linking node with its neighborhood. Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
We propose to use Binding the operation $\otimes$ to realize the function $\phi(\cdot)$ because it offers a means to balance the discriminative power of the resulting model against its generalization ability. We illustrate the preceding claim with an example. Let $v\in\mathcal{V}$ and $w\in\mathcal{V}$ be two nodes in a graph. Suppose $\phi(\cdot)$ is realized using the Binding Operation $\otimes$ and $\psi_{k}(\cdot)$ is realized using the Bundling Operation. Considering only 1-hop and 2-hop neighbors of $v$, the latent representation of $v$ and $w$ will be
$\begin{split}\mathbf{z}_{v}&=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}&=\mathbf{r}_{w}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(w)}\mathbf{r}_{m}\Big{)}\end{split}$ (1)
@@
It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired. On the other hand, if the encodings of the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome. When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater. The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
@@
Conversely, it is easy to see that Bundling is not a good choice for realizing $\phi(\cdot)$. Bundling of the node vector $\mathbf{r}_{v}$ with neighborhood information vector $\oplus_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$, a vector that has already been bundled with $k$-hop neighbors, will result in a vector that predominantly represents information from a node’s neighborhood, while leaving out essential information about the itself. This is due to the fact that Bundling is a bitwise majority operation; and hence, if bundling is used to realize $\phi(\cdot)$, the information from a node’s neighborhood suppresses the information about the node itself.
Table 1: Dataset Summary.
Dataset Nodes Edges Features Feature Type Classes Label Rate Edge Density
CORA 2708 5429 1433 $\{0,1\}$ 7 0.052 0.0004
CiteSeer 3327 4732 3703 $\{0,1\}$ 6 0.036 0.0004
PubMed 19717 44338 500 $(0,1)$ 3 0.003 0.0001
BlogCatalog 5196 171743 8189 $\mathbb{Z}^{+}$ 6 0.023 0.0127
Coauthor CS 18333 81894 6805 $\mathbb{Z}^{+}$ 15 0.016 0.0001
Coauthor Physics 34493 247962 8415 $\mathbb{Z}^{+}$ 5 0.002 0.0001
DBLP 17725 105781 768 $\mathbb{R}$ 4 0.004 0.0006
@@
Improving injectivity via Rotation. Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model. For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq. (1). Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$. Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as $\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}$. Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different. This poses a problem is terms of discriminative power of the model, leading to a lack of expressive power in our HD-algorithm. To address this issue, we utilize the Rotation $\Pi$ operator as it protects/preserves information. To be more specific, with the incorporation of the rotation/permutation operator, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes to
@@
where $\Pi(\cdot)$ and $\Pi\;\Pi(\cdot)$ denote single and double rotation operations respectively. The preceding equation ensures that $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since rotation of a HD vector yields a HD vector that is dissimilar/orthogonal to the original. That is,
@@
$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$ (2)
4.3 Node Label Prediction using HDGL

### v3

@@ 3.3. Graph Neural Networks
Representational Power of GNN. The representational power of GNNs can be characterized in terms of their ability to distinguish between different graph structures (Xu et al., 2019). The greater the representational power, the more minute the differences between graphs that can be detected by the GNN. GNN have been shown to be at most as powerful as the Weisfeiler-Lehman (WL) graph isomorphism test, which iteratively updates node features by aggregating neighbors’ features using an injective aggregation function (Weisfeiler and Leman, 1968; Huang and Villar, 2021). The injectivity of the aggregation function ensures that the differences in node neighborhoods result in differences in the resulting aggregations (Huang and Villar, 2021). Consequently, GNN are at most as powerful as the WL test in distinguishing graph structures if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme: $\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$ where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
4. Hyper-Dimensional Graph Learner
HDGL constructs a HD representation of graphs that matches the expressive power of GNN. HDGL exploits the properties of HD operations, namely, Bundling ($\oplus$), Binding ($\otimes$), and Rotation ($\Pi$) to ensure that graphs with similar node features and topologies map to similar HD representations. Specifically, HDGL: (i). Maps node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$. (ii). Constructs a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$. (iii). Uses $\mathbf{z}_{v}$ to predict node labels and links in the graph. We proceed to describe each of these steps in detail.
Figure 1. Mapping node features to a binary space using Random Hyperplanes: points are assigned 1 or 0 based on their position relative to each hyperplane.
[…]
4.1. Mapping Node Features to HD-space
@@
The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$. We do this using random hyperplane tessellations (Dirksen et al., 2022), a type of Locality Sensitive Hashing (Indyk and Motwani, 1998; Charikar, 2002). The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane. With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node. More precisely, let $\mathbf{Q}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{q}_{1}^{T},\cdots\mathbf{q}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$. For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$. Using $\mathbf{Q}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{Q}\mathbf{x}_{v}+\Gamma)$. The offset $\Gamma$ when it is non-zero ensures that the sampled hyperplane does not pass through the origin. Note that the resulting binary sketch of each node is in fact a $\beta$-dimensional binary HD vector. Nodes with similar features will have similar HD representations; and nodes with dissimilar features will have dissimilar HD representations.
4.2. Neighborhood Aggregation Scheme
@@
To simplify the mapping of nodes to their HD encodings, we assume that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have identical local topologies. Under this assumption, we can approximate injectivity of the embedding $\mathbf{z}_{v}$ of each node $v$ to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows: $\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}$ where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
Ensuring that similar nodes have similar HD representations. Suppose for a moment that the Binding operation $\otimes$ is used to realize $\varphi_{k}(\cdot)$. Because binding of two similar vectors results in dissimilar vectors, it would adversely impact the utility of the HD representation for predicting node labels based on node features and local node topologies.
@@
In contrast, we use Bundling operation because it is permutation-invariant and is well-suited for implementing the functions $\varphi_{k}(\cdot)$ under the assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have identical local topologies. More precisely, $\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes bundling the HD representations of all of the $k$-hop neighbors of $v$. This operation is analogous to finding a high-dimensional vector that is most similar to all of the input high-dimensional vectors, similar to finding the mean in Euclidean spaces and like the mean, it is not necessarily injective. However, as noted by (Xu et al., 2019), GNNs that use the mean to aggregate features over node neighborhoods (Kipf and Welling, 2017; Hamilton et al., 2017) perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is typically the case in real-world graphs.
Binding for linking node with its neighbors. Having chosen Bundling to realize $\varphi_{k}(\cdot)$, we proceed to consider how to realize $\phi(\cdot)$, which has to couple the representation of a node obtained using bundling (see above) with the representations of its neighbors.
As we will see shortly, if we choose to realize $\phi(\cdot)$ using the Binding operation $\otimes$, we obtain a representation that strikes a reasonable balance between discriminating between dissimilar graphs and generalizing over similar graphs. Let $v\in\mathcal{V}$ and $w\in\mathcal{V}$ be two nodes in a graph. Suppose $\phi(\cdot)$ is realized using the Binding Operation $\otimes$ and $\psi_{k}(\cdot)$ is realized using the Bundling Operation. Considering only 1-hop and 2-hop neighbors of $v$, the latent representation of $v$ and $w$ will be
(1) $\begin{split}\mathbf{z}_{v}&=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}&=\mathbf{r}_{w}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(w)}\mathbf{r}_{m}\Big{)}\end{split}$
@@
It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired. On the other hand, if the encodings of the features of the 2-hop neighbors of both $v$ and $w$ differ, $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar as desired. When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater. The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
@@
On the other hand, it is easy to see that Bundling is not a good choice for realizing $\phi(\cdot)$. Because bundling is a bitwise majority operation, bundling the node vector $\mathbf{r}_{v}$ with neighborhood information vector $\oplus_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$; results in a vector where the information from a node’s neighborhood suppresses the information about the node’s features, which is an undesirable property.
@@
Improving injectivity via Rotation. Because Binding is associative and commutative, it can adversely impact the injectivity of $\varphi_{k}(\cdot)$ and hence the representational power of our HD model. For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq. (1). Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$. Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as $\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}$. Since, Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different. This adversely impacts the representation’s ability to discriminate between dissimilar graphs. To address this issue, we utilize the Rotation $\Pi$ operator. Now, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes:
@@
where $\Pi(\cdot)$ and $\Pi\;\Pi(\cdot)$ denote single and double rotation operations respectively. Because the rotation of HD vector yields a HD vector that is dissimilar or orthogonal to the original, now $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar. That is,
@@
(2) $\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$
Input: $\mathcal{G}=(\mathcal{V},\mathcal{E})$, Node Features $\mathbf{x}_{v}$, Train Nodes $\mathcal{T}_{\ell}\subset\mathcal{V}$ for each class $\forall\ell\in\{1,\cdots,L\}$ and Unlabeled Test Nodes $\mathcal{V}_{\text{test}}$
Sample $\mathbf{Q}:[\mathbf{q}_{1}^{T}\sim\mathcal{N}(0,I_{d}),\cdots,\mathbf{q}_{\beta}^{T}\sim\mathcal{N}(0,I_{d})]$;
Sample $\Gamma\sim[-\lambda,\lambda]^{\beta}$;
for $v\in\mathcal{V}$ do
       $\mathbf{r}_{v}\leftarrow\text{sign}(\mathbf{Q}\mathbf{x}_{v}+\Gamma)$;
for $v\in\mathcal{V}$ do
       $\mathbf{z}_{v}\leftarrow\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}$;
if Node Label Prediction Task then
      for $\ell\in\{1,\cdots,L\}$ do
             $\mathbf{c}_{\ell}\leftarrow\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i};\quad$Computing Label Hyper Vectors;
      for $u\in\mathcal{V}_{\text{test}}$ do
             $\hat{y}_{u}\leftarrow\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}];\>$Test-set Inference;
if Link Prediction Task then
       $\mathbf{e}^{+}\leftarrow\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v})$ and $\mathbf{e}^{-}\leftarrow\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v})$;
       Compute $\mathbf{Z}^{+}$ and $\mathbf{Z}^{-}$ using Eqn. (3);
       $\mathbf{D}^{+}\leftarrow dist(\mathbf{Z}^{+},\mathbf{Z})$ and $\mathbf{D}^{-}\leftarrow dist(\mathbf{Z}^{-},\mathbf{Z})$;
       for $i=1$ to $N$  do
             for $j=1$ to $N$ do
                   if $D^{+}_{ij}<D^{-}_{ij}$ then
                         $\hat{A}_{ij}\leftarrow\sigma\big{(}(1-D^{+}_{ij})+D^{-}_{ij}\big{)}$;
                  else
                         $\hat{A}_{ij}\leftarrow\sigma\big{(}D^{+}_{ij}-(1-D^{-}_{ij})\big{)}$;
Algorithm 1 Pseudo-code for HDGL
4.3. Node Label Prediction using HDGL